# Iterative Repair with Weak Verifiers for Few-shot Transfer in KBQA with Unanswerability

### v2

@@ 5 Experiments
Model WebQSP $\rightarrow$ GrailQAbility WebQSP $\rightarrow$ GraphQAbility
Overall Answerable Unanswerable Overall Answerable Unanswerable
@@
Table 1: Performance of different models on two datasets for few-shot KBQA transfer with unanswerability.
We now present experimental evaluation of FUn-FuSIC. For few-shot KBQA transfer with unanswerability, we address the following research questions. (R1) How does FUn-FuSIC compare against SoTA models for KBQA, suitably adapted for this setting? (R2) How does FUn-FuSIC perform across different categories of unanswerability? (R3) How do the different components of FUn-FuSIC contribute to its performance? In addition, for answerable KBQA few-shot transfer, we ask: (R4) How does FUn-FuSIC compare against SoTA KBQA models for this setting?
@@ 5.1 Experimental Setup
For few-shot KBQA transfer, available datasets have only answerable questions Patidar et al. (2024), while our task needs the target dataset to contain unanswerable questions as well. So we construct our own transfer datasets, augmenting existing ones. GrailQAbility is the only available KBQA dataset with unanswerable questions Patidar et al. (2023). This was carefully designed starting from GrailQA Gu et al. (2021), which has only answerable questions, by systematically deleting schema and data elements from the back-end KB to introduce different categories of unanswerability into the queries. We use this as one of our targets.
GraphQA Su et al. (2016) is another popular KBQA dataset. This has the same back-end KB (Freebase) as GrailQA. We create our second target dataset using GraphQA, by replacing its KB with the modified KB in GrailQAbility. This introduces unanswerability into GraphQA questions and we label these appropriately as schema-level or data-level unanswerable. We name this dataset GraphQAbility. We create the two test sets by selecting 250 answerable and 250 unanswerable questions uniformly at random from their test sets. We create few shots by selecting 100 questions (50 answerable and 50 unanswerable) uniformly at random from the dev set and train set respectively.
@@
WebQSP training set has 2,858 real user questions, which are manually annotated with logical forms. This is quite different from GraphQA and GrailQA which contain algorithmically generated logical forms, verbalized by crowd-workers. Because of this difference in nature of questions, these are challenging transfer datasets, beyond the unaswerability gap.
Model WebQSP $\rightarrow$ GrailQAbility WebQSP $\rightarrow$ GraphQAbility
Schema Level Data Level Schema Level Data Level
@@
As few-shot transfer for KBQA with unanswerability is a novel task, there are no existing baselines. For in-domain KBQA with unanswerability, RetinaQA Faldu et al. (2024) and Pangu Gu et al. (2023). which has been adapted for unanswerable questions, are the two SoTA models. For these, we use the available code. [^2] https://github.com/dair-iitd/RetinaQA, [^3] https://github.com/dki-lab/Pangu More details are in Appendix A.10.
FuSIC-KBQA is the SoTA model for few-transfer for KBQA with only answerable questions. Instead of retrieve-then-generate, KB-Binder Li et al. (2023) follows a generate-then-ground approach. It is the SoTA for in-domain few-shot KBQA, and overall, FuSIC-KBQA and KB-Binder outperform all other supervised and LLM-equipped KBQA models adapted for few-shot transfer Patidar et al. (2024). We use available code for KB-Binder [^4] https://github.com/ltl3A87/KB-BINDER, and our own implementation for FuSIC-KBQA since source code is not available. To adapt these two baselines for unanswerability, for fair comparison, we modify their logical form generation prompt in the same fashion as PUn for FUn-FuSIC. Additionally, for FuSIC-KBQA, we remove execution-guided feedback (EGF) since it fails for unanswerability. We denote this model FuSIC-KBQA-U. Observe that FuSIC-KBQA-U can also be seen as an ablation of FUn-FuSIC, without FUn. More details about use of KB-Binder are in the Appendix A.12.
We use $\mathcal{L}=$gpt-4-0613 for all three models. For a fair comparison, we provide all three LLM-equipped models the same aggregated prompt limit for a question. To achieve this, to FUn-FuSIC we provide zero-shot generation and $n=4$ FUn iterations, to FuSIC-KBQA-U 5-shot generation, and to KB-Binder 25-shot generation.
Though FUn-FuSIC and FuSIC-KBQA allow flexible use of multiple supervised retrievers, for meaningful comparison with RetinaQA, we adapt RetinaQA as retriever for FUn-FuSIC and FuSIC-KBQA-U. More details about KB-Binder and FuSIC-KBQA’s generation is in Appendix A.12, and details of FuSIC-KBQA’s retriever along compute infrastructure are in Sec. A.9.
@@
Since our primary task is generation of logical forms, our primary focus in evaluation is on logical forms as well. For logical form evaluation, the existing exact match measure (EM) Ye et al. (2022) works only for s-expression as the language. Since FuSIC-KBQA-U and FUn-FuSIC output logical forms in SPARQL, and Pangu, RetinaQA and KB-Binder in s-expression, we propose a new measure that works for both languages. This measure, EM-s, considers two logical forms to be equivalent if these contain identical sets of relations and entities, and additionally return identical answers upon execution. This is a necessary but not sufficient condition for logical form equivalence. So, we compared EM-s with EM in settings where both are applicable, and found $>98\%$ agreement. More details are in the Appendix A.3.
Since logical form evaluation is not completely accurate (logical forms different from the gold one may still be correct), following existing literature Patidar et al. (2023); Faldu et al. (2024), we also evaluate answers using regular F1 (F1(R)), also lenient F1 (F1(L)), which does not penalize the original answer for the complete KB.
Answerable Unanswerable
F1 EM-s F1(L) F1(R) EM-s
FUn-FuSIC 57.0 46.0 92.0 90.0 62.0
scUn $\Rightarrow$ sc 63.0 52.0 66.0 64.0 36.0
w/o syntax 57.0 46.0 92.0 90.0 62.0
w/o kb-inc 51.7 42.0 92.0 88.0 44.0
w/o q-lf 47.7 38.0 55.8 48.0 10.0
w/o ans-inc 55.0 44.0 92.0 90.0 62.0
Table 3: Ablation performance of FUn-FuSIC (removing individual components with replacement) on subset of WebQSP $\rightarrow$ GraphQAbility. scUn $\Rightarrow$ sc denotes replacing scUn with self consistency. Other rows remove verifiers for syntax error (w/o syntax) (V1), KB inconsistency (w/o kb-inc) (V2), question logical form disagreement (w/o q-lf) (V3) and answer incompatibility (w/o ans-inc) (V4).
@@ 5.2 Unanswerability Setting
We first address research question R1. Performances of the different models for the few-shot transfer setting with unanswerability are recorded in Tab. 1. Note that the Overall columns determine superiority of one model over another, while the Answerable and Unanswerable columns provide further drill-down for analysis.
First, we observe that FUn-FuSIC significantly outperforms all baselines in terms of EM-s, and performs at par with FuSIC-KBQA-U and significantly better than all other models in F1. However, the other LLM-equipped models do not far surpass supervised models. In fact, all 4 models perform almost at par for GraphQAbility, and KB-Binder performs worse than the other 3 for GrailQAbility. This establishes usefulness of FUn+scUN for few-shot transfer KBQA with unanswerability. Secondly, each model trades off performance differently between answerable and unanswerable questions. RetinaQA, Pangu and also KB-Binder fare better for unanswerable questions, while FUn-FuSIC and FuSIC-KBQA-U fare better for answerable ones. However, FUn-FuSIC achieves the best balance between the two.
We next briefly address research question R2. Performance of different models for different categories of unanswerability are recorded in Tab. 2. Models struggle to fare well across both data-level and schema-level unanswerability. FuSIC-KBQA-U performs the best for data-level while performing poorly (in terms of EM-s) for schema-level. Conversely, RetinaQA performs well for schema-level, but has poor data-level EM-s. FUn-FuSIC performs poorer for data-level compared to schema-level but achieves the best balance by far between the two categories among all models.
We next address research question R3. Tab. 3 records performance for different ablations of FUn-FuSIC. Note that this experiment is performed on a subset of the test data (50 questions drawn randomly from the each of the answerable and unanswerable categories). The biggest benefit, for both answerable and unanswerable questions, comes from Question - Logical Form Disagreement verifier. KB Inconsistency and Answer Incompatibility verifiers also make significant contributions to the performance. This demonstrates the usefulness of weak verifiers. Replacing scUn with self-consistency, as expected, leads to a drastic drop in unanswerable performance (though this comes with a benefit for answerable questions).
Model WebQSP $\rightarrow$ WebQSP $\rightarrow$
GrailQA-Tech GraphQA-Pop
@@ 5.3 Answerable Setting
We now address research question R4 (answerable-only KBQA). We use two datasets from existing literature Patidar et al. (2024), including the hardest one (WebQSP $\rightarrow$ GraphQA-Pop). [^5] https://github.com/dair-iitd/FuSIC-KBQA/ For consistency with existing literature, here all models use TIARA Shu et al. (2022) as the retriever.
This setting admits two sub-settings: (A) the models have knowledge that all questions are answerable, and (B) though all questions are answerable, the models do not have this prior knowledge.
Setting (A) has been studied for KBQA Patidar et al. (2024), and FuSIC-KBQA is the established SoTA model, outperforming a host of supervised and LLM-based models adapted for the task. In this setting, FUn-FuSIC requires three simple modifications. (i) PUn is replaced with prompt for answerability, (ii) In FUn, category of V4b (empty answer) changes from weak verifier to strong verifier, and (iii) scUn is replaced by self-consistency.
The first two rows in Tab. 4 record performance for setting (A). FUn-FuSIC significantly outperforms FuSIC-KBQA on both datasets, creating a new SoTA for this setting. This shows the usefulness of iterative repair with a suite of strong and weak verifiers followed by self-consistency for KBQA transfer, even without unanswerability.
In the realistic second setting (B) – not studied before – the models make predictions assuming unanswerability. Here, we evaluate FuSIC-KBQA-U and FUn-FuSIC as in Sec. 5.2. The bottom two rows of Tab. 4 record their performance. We see that FUn-FuSIC outperforms FuSIC-KBQA by a very large margin. This further establishes the usefulness of scUn when guarantees about answerability are not available.
@@ 5.4 Error Analysis
$l^{*}\neq$ NK, ${\hat{l}}\neq$NK, $l^{*}\neq{\hat{l}}$ 9.8
Table 5: FUn error analysis on WebQSP $\rightarrow$ GraphQAbility. $l^{*}$ ${\hat{l}}$ denote gold & generated logical forms. Retrieval error means retrieval $r$ is missing $\geq 1$ KB elements (class, relation, entity) necessary for $l^{*}$. Generation error implies ${\hat{l}}\neq l^{*}$ despite correct retrieval.
For WebQSP $\rightarrow$ GraphQAbility, we analyzed questions whose logical forms are judged incorrect (EM-s $<1$). See results in Tab. 5. We found three main causes for generation errors. (1) Some questions are inherently ambiguous, admitting multiple valid logical forms $l_{1}$ and $l_{2}$ in the original complete KB, though only one is recognized as the gold ($l^{*}=l_{1}$). Deletion to introduce unanswerability eliminates $l_{1}$, so we update $l^{*}=$ NK, and the prediction ${\hat{l}}=l_{2}$ is unfairly penalized. (2) $l^{*}=l_{1}$, ${\hat{l}}=l_{2}$, where $l_{1}\neq l_{2}$ while being semantically equivalent, but are incorrectly judged non-equivalent by EM-s. (3) FUn is genuinely unable to generate $l^{*}$ or any semantic equivalent within its iteration limit.
We also estimated the accuracy of specific weak verifiers by manual analysis of $\sim$100 instances from test. Verifier V3 for Question - Logical Form Disagreement has accuracy of 93.6% overall, with its Query to Natural Language Back-translation component having 90%. Verifier V4b for Empty Answer has 67.8% (this is expected since $\sim$25of questions have ($l^{*}\neq$ NK, $A^{*}=$ NA).
@@ 6 Conclusions
For real-world robust, low-resource KBQA, we have proposed the novel task of few-shot transfer learning with unanswerability. We have proposed FUn-FuSIC that augments for unanswerability the SoTA few-shot answerable-only KBQA transfer model by (i) iterative repair using feedback from a suite a strong and weak verifiers – including a novel back-translation based verifier – to create a set of candidate logical forms, and (ii) assessing this candidate set to detect unanswerability (and its category) and/or identify the best logical form using self consistency adapted for unanswerability.
Using two newly created datasets for this novel task, we show that FUn-FuSIC outperforms adaptations of SoTA models for this setting, and also for answerable few-shot transfer KBQA. Error analysis suggests that performing well across categories of unanswerability for few-shot transfer is still a challenge and should be a focus of further research.
7 Limitations
Since LLM inference involves randomness, ideally experiments should be repeated for multiple runs and results should report averages and error bars. Unfortunately, we were not able to do this due to the prohibitive cost of GPT-4, and our results are based on single runs.
While GPT-4 is currently the best performing LLM, it is proprietary as well as expensive. Ideally, evaluation should involve open-source freely accessible LLMs as well. We expect performance of all LLM-based approaches to drop when GPT-4 is replaced by an open LLM. Earlier research has shown that models using Mistral are still able to outperform fully supervised models for answerable few-shot transfer Patidar et al. (2024). Whether this trend holds for the unanswerable setting is an open question. That said, following current trends, we expect the ability of open LLMs to steadily improve in the coming years.
8 Risks
At the highest level, our work reduces risk compared to existing KBQA systems, which when improperly adapted in a low-resource setting, incorrectly answer unanswerable questions, without admitting lack of knowledge. However, can incorrectly inferring unanswerability, citing lack of knowledge when the knowledge is in fact available, generate a new category of risk? While we cannot imagine such a risk at the present time, this may require more thought. In any case, KBQA models for unanswerability should strive to minimize this type of error, along with the other types.

### v3

@@ 5 Experiments
WebQSP $\rightarrow$ GrailQAbility WebQSP $\rightarrow$ GraphQAbility
Model Overall Answerable Unanswerable Overall Answerable Unanswerable
@@
Table 1: Performance of different models on two datasets for few-shot KBQA transfer with unanswerability. Answerable and Unanswerable record performance for corresponding subsets and Overall for the entire dataset.
We now present experimental evaluation of FUn-FuSIC. First, for few-shot KBQA transfer with unanswerability, we address the following research questions. (R1) How does FUn-FuSIC compare against SoTA KBQA models suitably adapted for this setting? (R2) How does FUn-FuSIC perform across different categories of unanswerability? (R3) How do the different components of FUn-FuSIC contribute to its performance? Then, for answerable KBQA few-shot transfer, we ask: (R4) How does FUn-FuSIC compare against SoTA KBQA models for this setting?
@@ 5.1 Experimental Setup
For in-domain and answerable KBQA, the three most popular datasets are GrailQA Gu et al. (2021), GraphQA Su et al. (2016) and WebQSP Yih et al. (2016). All of these have the same back-end KB (Freebase). For few-shot KBQA transfer, the only available datasets also have only answerable questions Patidar et al. (2024). GrailQAbility is the only available KBQA dataset with unanswerable questions Patidar et al. (2023). This was constructed starting from GrailQA Gu et al. (2021) by systematically deleting schema and data elements from the back-end KB to introduce different categories of unanswerability into the queries.
@@
[…]
WebQSP $\rightarrow$ GrailQAbility WebQSP $\rightarrow$ GraphQAbility
Model Schema Level Data Level Schema Level Data Level
@@
As few-shot transfer for KBQA with unanswerability is a novel task, there are no existing baselines. For in-domain KBQA with unanswerability, RetinaQA Faldu et al. (2024) and the unanswerability-adapted version of Pangu Gu et al. (2023) are the SoTA models. For these, we use the available code. [^2] https://github.com/dair-iitd/RetinaQA, [^3] https://github.com/dki-lab/Pangu More details are in the Appendix (Sec. A.8.3).
FuSIC-KBQA is the SoTA model for few-transfer for KBQA with only answerable questions. KB-Binder Li et al. (2023) is the SoTA for in-domain few-shot KBQA. Overall, FuSIC-KBQA and KB-Binder outperform all other supervised and LLM-equipped KBQA models adapted for few-shot transfer Patidar et al. (2024). We use available code for KB-Binder [^4] https://github.com/ltl3A87/KB-BINDER, and our own implementation for FuSIC-KBQA. To adapt these two baselines for unanswerability, for fair comparison, we modify their logical form generation prompt in the same fashion as PUn for FUn-FuSIC. Additionally, for FuSIC-KBQA, we remove execution-guided feedback (EGF) since it fails for unanswerability. We denote this model FuSIC-KBQA-U. Observe that FuSIC-KBQA-U can also be seen as an ablation of FUn-FuSIC, without FUn. More details about KB-Binder and FuSIC-KBQA are in the Appendix (Sec. A.9.1).
We use $\mathcal{L}=$gpt-4-0613 for all LLM-equipped models. For fair comparison, we allocate to all such models the same maximum aggregated prompt length for a question. This is satisfied by equipping FUn-FuSIC with zero-shot generation and $n=4$ FUn iterations, FuSIC-KBQA-U with 5-shot generation and KB-Binder with 25-shot generation.
Though FUn-FuSIC and FuSIC-KBQA allow flexible use of multiple supervised retrievers, for meaningful comparison with RetinaQA, we adapt RetinaQA as retriever for FUn-FuSIC and FuSIC-KBQA-U. More details about FuSIC-KBQA’s retriever and compute infrastructure are in the Appendix (Sec. A.8.2).
@@
For KBQA as semantic parsing task, evaluation of logical forms is primary. For this, the existing EM measure Ye et al. (2022) is defined only for logical forms represented using s-expressions. FuSIC-KBQA-U and FUn-FuSIC output logical forms in SPARQL, and Pangu, RetinaQA and KB-Binder in s-expression. So we propose a new measure EM-s that checks approximate equivalence for a pair of programs either in SPARQL or s-expression. More details are in the Appendix (Sec. A.5).
As in standard KBQA evaluation, we also evaluate answers. This is a secondary evaluation for giving the benefit of the doubt for getting the right answer, possibly via a logical form not equivalent to the gold-standard according to EM-s. For answer evaluation in KBQA with unanswerability,  Patidar et al. (2023) introduced lenient F1, denoted F1(L), in addition to regular F1, denoted F1(R). F1(L) relaxes F1(R) by not penalize the original answer for the complete KB. Note that obtaining the right answer by chance has much higher probability than for logical forms, particularly for unanswerable questions with NA as the correct answer.
WebQSP $\rightarrow$ GrailQAbility WebQSP $\rightarrow$ GraphQAbility
Model Answerable Schema L. UnAnswerable Data L. UnAnswerable Answerable Schema Level UnAns Data Level UnAns
F1 EM-s F1(L) F1(R) EM-s F1(L) F1(R) EM-s F1(L) EM-s F1(L) F1(R) EM-s F1(L) F1(R) EM-s
FUn-FuSIC 74.0 70.0 90.9 87.9 75.8 64.7 64.7 11.8 59.0 48.0 97.1 91.4 80.0 86.7 86.7 26.3
scUn $\Rightarrow$ sc 74.0 70.0 73.0 72.7 33.3 58.8 52.9 23.5 64.0 54.0 74.4 68.6 22.9 80.0 80.0 13.3
w/o syntax 67.3 64.0 87.9 90.9 42.4 76.5 70.6 17.7 57.0 46.0 81.0 76.9 26.9 75.8 75.0 16.7
w/o kb-inc 70.3 66.0 90.9 90.9 9.1 76.5 70.6 29.4 51.7 42.0 100.0 100.0 4.2 79.5 75.0 20.8
w/o q-lf 71.0 68.0 69.7 63.6 0.0 41.2 35.3 5.9 47.7 38.0 69.5 65.4 07.7 67.0 62.5 20.8
w/o ans-inc 72.0 70.0 85.9 84.9 33.3 70.6 70.6 23.5 55.0 44.0 80.8 76.9 26.9 79.5 79.2 25.0
Table 3: Ablation performance of FUn-FuSIC (removing individual components with replacement) on subset of WebQSP $\rightarrow$ GraphQAbility. scUn $\Rightarrow$ sc denotes replacing scUn with self consistency. Other rows remove verifiers for syntax error (w/o syntax) (V1), KB inconsistency (w/o kb-inc) (V2), question logical form disagreement (w/o q-lf) (V3) and answer incompatibility (w/o ans-inc) (V4). Evaluations are on 100 instances from test sets (50 answerable and 50 unanswerable questions sampled uniformly at random)
@@ 5.2 Unanswerability Setting
We first address research question R1. Performances of different models for few-shot transfer with unanswerability are recorded in Tab. 1. First, we observe that FUn-FuSIC significantly outperforms all baselines in terms of EM-s, and performs at par with FuSIC-KBQA-U and significantly better than all other models in F1. The other LLM-equipped models are not significantly better than the supervised models. All the baselines perform almost at par for GraphQAbility, and KB-Binder performs worse than the other 3 for GrailQAbility. This establishes usefulness of FUn equipped with scUn for few-shot transfer KBQA with unanswerability, beyond LLM-usage. Secondly, each model trades off performance differently between answerable and unanswerable questions. RetinaQA, Pangu and also KB-Binder fare better for unanswerable questions, while FUn-FuSIC and FuSIC-KBQA-U fare better for answerable ones. However, FUn-FuSIC achieves the best balance across the two subsets.
We next briefly address research question R2. Performance of different models for different categories of unanswerability are recorded in Tab. 2. All models struggle to fare well simultaneously for data-level and schema-level unanswerability. FuSIC-KBQA-U performs the best for data-level while performing poorly (in terms of EM-s) for schema-level. Conversely, RetinaQA performs well for schema-level, but has poor data-level EM-s. FUn-FuSIC outperforms other models in schema-level unanswerability while being slightly worse in data-level unanswerability. But among all models, it achieves the best tradeoff by far across unanswerability categories.
We next address research question R3. Tab. 3 records the ablation analysis of FUn-FuSIC. We see that the KB-Inconsistency verifier (V2) and the Q-LF Disagreement verifier (V3) lead to significant improvements in EM-s. The biggest benefit comes from V3 for both answerable and unanswerable questions. Without V3, the correct LF for schema-level unanswerability is almost never generated, though answer accuracy stays high, indicating inability to reason. Similarly, removing V2 reduces EM-s for schema-level unanswerable questions, with answer accuracy remaining high, indicating that answers are often correct despite flawed logical forms. The Answer Incompatibility verifiers (V4) also makes significant contributions to the performance. This analysis highlights the necessity of a mix of weak and strong verifiers for structural and semantic validity. Beyond verifiers, replacing scUn with self-consistency, as expected, leads to a drastic drop in unanswerable performance (though this comes with a benefit for answerable questions).
WebQSP $\rightarrow$ WebQSP $\rightarrow$
Model GrailQA-Tech GraphQA-Pop
@@ 5.3 Answerable Setting
We now address research question R4 for answerable-only KBQA transfer. We use two datasets from existing literature Patidar et al. (2024), including the hardest one (WebQSP $\rightarrow$ GraphQA-Pop). [^5] https://github.com/dair-iitd/FuSIC-KBQA/ For enabling comparison with earlier results, we use TIARA Shu et al. (2022) as the retriever for all models in this experiment.
This setting admits two sub-cases: (A) the models have knowledge that all questions are answerable, and (B) though all questions are answerable, the models do not have this knowledge.
Setting (A) has been studied for KBQA Patidar et al. (2024), and FuSIC-KBQA is the established SoTA model, outperforming a host of supervised and LLM-based models adapted for the task. To adapt for this setting, FUn-FuSIC requires three simplifications. (i) PUn is replaced with prompt for answerability, (ii) In FUn, V4b (empty answer) is moved from the set of weak verifiers to that of strong verifier, and (iii) scUn is replaced by standard self-consistency.
The first two rows in Tab. 4 record performance for setting (A). FUn-FuSIC significantly outperforms FuSIC-KBQA on both datasets, creating a new SoTA for this setting. This shows the usefulness of iterative repair with a suite of strong and weak verifiers followed by self-consistency for few-shot KBQA transfer, even without unanswerability.
In the more realistic setting (B), which has not been studied before, the models make predictions assuming unanswerability. Here, we evaluate FuSIC-KBQA-U and FUn-FuSIC as in Sec. 5.2, only there are no truly unanswerable questions. The bottom two rows of Tab. 4 record the performance of the two models in this setting. We see that FUn-FuSIC outperforms FuSIC-KBQA by a very large margin. This further establishes the usefulness of scUn when guarantees about answerability are not available.
@@ 5.4 Error Analysis
$l^{*}\neq$ NK, $^{\hat{l}}\neq$NK, $l^{*}\neq{\hat{l}}$ 9.8
Table 5: FUn error analysis on WebQSP $\rightarrow$ GraphQAbility. $l^{*}$ $^{\hat{l}}$ denote gold & generated logical forms. Retrieval error means retrieval $r$ is missing 1$\geq$ 1≥ 1 KB elements (class, relation, entity) necessary for $l^{*}$. Generation error implies ${\hat{l}}\neq l^{*}$ despite correct retrieval.
For WebQSP $\rightarrow$ GraphQAbility, we analyzed questions for which logical forms generated by FUn-FuSIC are incorrect (EM-s <1<1< 1). Results are in Tab. 5. We found three main causes for generation errors. (1) Some questions are inherently ambiguous, admitting multiple valid logical forms $l_{1}$ and $l_{2}$ in the original complete KB, though only one is recognized as the gold ($l^{*}=l_{1}$). Deletion to introduce unanswerability eliminates $l_{1}$, so that that $l^{*}=$ NK, and the prediction ${\hat{l}}=l_{2}$ is unfairly penalized. (2) Here, $l^{*}=l_{1}$ and the prediction ${\hat{l}}=l_{2}$, such that $l_{1}\neq l_{2}$ but are semantically equivalent. $l_{1}$ and $l_{2}$ are incorrectly judged non-equivalent by EM-s. (3) Here, FUn is unable to generate $l^{*}$ or any semantic equivalent of it within its iteration limit.
@@ 6 Conclusions
For real-world robust and low-resource KBQA, we have proposed the novel task of few-shot transfer learning with unanswerability. We have introduced a new notion (FUn) of iterative feedback guided repair for answerable as well as unanswerable questions. FUn (i) uses feedback from a diverse suite a strong and weak verifiers – including a novel back-translation based verifier – to create a set of candidate logical forms, and (ii) assesses this candidate set to either to detect unanswerability (and its category) or identify the best logical form using self consistency adapted for unanswerability (scUn). We propose FUn-FuSIC that replaces the existing the iterative strategy, that assumes answerability of questions, with FUn in the SoTA few-shot answerable-only KBQA transfer model (FuSIC-KBQA). Using two newly created datasets for this novel task, we show that FUn-FuSIC significantly outperforms adaptations of FuSIC-KBQA and other SoTA models for this setting, and also for answerable few-shot transfer KBQA.
Our error analysis suggests that performing well across categories of unanswerability for few-shot transfer is still a challenge for KBQA and should be a focus of further research. We have made our datasets and other resources public  [^6] https://github.com/dair-iitd/FUn-FuSIC.
Limitations
Since LLM inference involves randomness, experiments should ideally be repeated for multiple runs and results should report averages and error bars. Unfortunately, we were not able to do this due to the prohibitive cost of GPT-4, and our results are based on single runs.
While GPT-4 is currently the best performing LLM, it is proprietary as well as expensive. Ideally, evaluation should include open-source freely accessible LLMs as well. We expect performance of all LLM-based approaches to drop when GPT-4 is replaced by a less powerful, open LLM. Nonetheless, earlier research has shown that models with Mistral instead of GPT-4 still outperform fully supervised models for answerable few-shot transfer Patidar et al. (2024). Whether this trend holds for the unanswerable setting is an open question. That said, following current trends, we expect the ability of open LLMs to steadily improve in the coming years.
Risks
At the highest level, our work reduces risk compared to existing KBQA systems, which when inadequately adapted in a low-resource setting, incorrectly answer unanswerable questions, without acknowledging lack of knowledge. However, can incorrectly inferring unanswerability, citing lack of knowledge when knowledge is in fact available, be a new type of risk? While we cannot imagine such a risk at the present time, this may require more careful consideration. In any case, KBQA models for unanswerability should strive to minimize this type of error, along with the other types.
Acknowledgments
Mausam is supported by a contract with TCS, grants from IBM, Verisk, Huawei, Wipro, and the Jai Gupta chair fellowship by IIT Delhi. Indrajit would like to thank KnowDis AI for supporting his participation in the conference. Riya would like to thank Graviton Research Capital for supporting her participation in the conference. The authors would like to thank the IIT-D HPC facility for its computational resources. We also thank Microsoft Accelerate Foundation Models Research (AFMR) program that provided us access to OpenAI models. We are also thankful to Mayur Patidar for helpful discussions.